A 4.1-kg block of ice originally at 263 K is placed in thermal contact with a 13.1-kg block of silver (cAg = 233 J/kg-K) which is initially at 1064 K. The H2O - silver system is insulated so no heat flows into or out of it.

1) At what temperature will the system achieve equilibrium?

________K
Answer #1


heat gained by ice = Mice*Cice*dT1 + mice*Lf + mice*cw*dT2


Qgain = (4.1*2100*(273-263)) + (4.1*334000)+(4.1*4190*(T-273))


Qgain = 1455500 + 17179T - 4689867

Qgain = 17179T - 3.234367e+6


heat lost by silver = mAg*cAg*dT3


Qloss = 13.2*233*(1064-T) = 3272438.4 - 3075.6T

from pricilple of calorimetry

Qlosss = Q gain

3272438.4 - 3075.6T = 17179T - 3.234367e+6


T = 321 K